OpenStudy (anonymous):

solve using multiplication principle -6x>1/13 the solution set is {x/x type an inequality symbol and a fraction

OpenStudy (saifoo.khan):

\[-6x > \frac{1}{13}\] \[-x > \frac{1}{78}\] \[x < - \frac{1}{78}\]